# aoas/sm_enkf

### v2

@@ 3 Convergence of Ensemble Kalman Filter
[…]
@@
Assumption 1 (i) and (ii) are two versions of the condition on the model operator $\mathcal{M}_{t}$. If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1. If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite interval. However, strongly nonlinear systems may not satisfy the Lipschitz condition. In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which is related to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$. Assumption 2 assumes the eigenvalues of the three basic covariances $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity. From Lemma S.3.1 of the SM, the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ satisfy the conditions for the thresholding and bandable(or circular bandable) covariance classes under Assumption 2 and Assumptions 1-2, respectively. Assumption 3 puts a mild condition on the observation operator $\mathbf{H}_{t}$.
The following proposition provides a basic result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$ is an estimator of the Kalman gain matrix assuming $\mathbf{R}_{t}$ is known.
Proposition 3.1.
Suppose Assumption 2 holds and $\lambda_{\min}(\mathbf{R}_{t})\geq\varepsilon_{0}$ for a positive $\varepsilon_{0}$,
@@
The last inequality is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$. Proposition 3.1 suggests that the upper bound of the estimation error in $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$. Proposition 3.1 holds under weak conditions and actually is valid for both fixed and high dimensions. Note that under fixed dimension, the sample covariance is consistent with the underlying population covariance at the $\sqrt{n}$ rate, that is $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$. The results for the high dimensions are shown later. In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown. Corollary S.3.1 of the SM provides an upper bound of the estimation error of the $\hat{\mathbf{K}}_{t}$ to $\mathbf{K}_{t}$ when both $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
@@
Based on the result in Proposition 3.1, we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.5)-(2.6) by firstly establishing the convergence of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ to $\boldsymbol{\Sigma}_{t}^{f}$. For the fixed dimension case, the sample covariance matrix is consistent such that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$. However, when the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.9) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ (Bai and Yin, 1993), which is a major threat to the validity of the EnKF method.
@@
Next, we present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and their covariance classes before we give the convergence of the EnKF.
@@
3.1 Consistent estimation of high dimensional $\boldsymbol{\Sigma}_{t}^{f}$
@@
The banding estimator (Bickel and Levina, 2008b) of the forecast error covariance matrix with bandwidth $k$ is $B_{k}(\mathbf{S}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|i-j|\leq k)\big{]}$, where $0\leq k<p$ and $\mathbf{S}_{t}^{f}=[s_{tij}^{f}]_{p\times p}$ is the sample covariance of the EnKF forecast ensembles. The covariance matrix class suitable for the banding estimator is the following bandable covariance class, $\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}$ for some positive $\alpha$, C and $\varepsilon_{0}$. This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger.
@@
The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero. Cai et al. (2010) updated the banding estimator with a tapering alteration which tapers down the level of the thresholding gradually
$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{(}w_{ij}(k)s_{tij}^{f}\big{)}_{p\times p},$
@@
where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\left\{(k-|i-j|)_{+}-\left(k/2-|i-j|\right)_{+}\right\},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.
@@
Figure 1: Bandable and circular bandable covariance matrices: the average energy function $h(k)$ in Panels (a) and (b) and their heatmaps in Panels (c) and (d), where $h(k)$ is the average of the squares of the $k^{th}$ sub-diagonal entries.
@@
In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or its poles, the state vectors are highly dependent for the variable components between the two ends of the state vector. Thus we propose a circular form of the bandable structure, $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$ Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariance matrices that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$, where $h(k)$ is the average of the squares of the $k^{th}$ subdiagonal entries of a covariance defined by $h(k):=\frac{1}{2(p-k)}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=\frac{1}{p-k}\sum_{l=1}^{p-k}\sigma_{ll+k}^{2}.$ Figures 1 (a)-(b) show that the $h(k)$ function of covariances in $\mathcal{U}_{b}$ is a $L$-type curve as the subdiagonal index $k$ increases, while it is an $U$-type curve for covariances in $\mathcal{U}_{bc}$. For covariances in the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, the proposed mid-banding forecast error covariance estimator with bandwidths $k_{1}$ and $k_{2}$ is $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})=[s_{tij}^{f}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))]$, where $0\leq k_{1}<(p-k_{2})\leq p$.
@@
The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart. However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class. In such situations, the thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008a) may be employed, where $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i\text{ and }\lambda_{\min}(\boldsymbol{\Sigma})\geq\varepsilon_{0}>0\big{\}},\text{ for }0\leq\gamma<1.$ The thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|s_{tij}^{f}|\geq s)\big{]}$.
@@
Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally IID Gaussian or sub-exponentially distributed as in Assumption S.3.1 of the SM.
@@
    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$

### v3

@@ 3 Convergence of Ensemble Kalman Filter
We first provide the upper bound of the estimation error of the estimated Kalman Gain matrix. Throughout the paper, $\|\cdot\|$ denotes the $L_{2}$ (spectral) norm of a matrix or the $L_{2}$ norm of a vector, while other matrix norms will be specified explicitly. For two positive sequences $a_{n}$ and $b_{n}$, $a_{n}\asymp b_{n}$, $a_{n}\ll b_{n}$ and $a_{n}\gg b_{n}$ mean that $a_{n}/b_{n}$ is bounded away from zero and infinity, $a_{n}/b_{n}\rightarrow 0$ and $b_{n}/a_{n}\rightarrow 0$ as $n\rightarrow\infty$, respectively. The following assumptions are needed in the analysis.
@@
Assumption 1 (i) and (ii) are two versions of the condition on the model operator $\mathcal{M}_{t}$. If model $\mathcal{M}_{t}$ is linear, it automatically satisfies both (i) and (ii) of Assumption 1. If $\mathcal{M}_{t}$ is nonlinear and Lipschitz continuous, $\mathcal{M}_{t}(\cdot)$ has at most linear growth at infinite intervals. However, strongly nonlinear systems may not satisfy the Lipschitz condition. In that case, we assume $\mathcal{M}_{t}$ satisfies condition (ii), which is related to the tangent linear hypothesis (Thepaut and Courtier, 1991; Bouttier and Courtier, 1999) in the 4D-Var, where $\mathcal{M}_{t}$ is determined in the linear approximation $\mathcal{M}_{t}(\mathbf{x}_{t}+\mathbf{h})\approx\mathcal{M}_{t}(\mathbf{x}_{t})+\dot{M}_{t}\mathbf{h}$ in the vicinity of $\mathbf{x}_{t}$ for a perturbation $\mathbf{h}$, and the operator $\dot{M}_{t}$ is called the differential or tangent linear function of $\mathcal{M}_{t}$ at point $\mathbf{x}_{t}$. Assumption 2 assumes the eigenvalues of the three basic covariances $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$ and $\boldsymbol{\Sigma}_{0}^{a}$ are bounded away from zero and infinity. From Lemma S.3.1 of the SM, the eigenvalues of $\boldsymbol{\Sigma}_{t}^{f}$ satisfy the conditions for the thresholding and bandable(or circular bandable) covariance classes under Assumption 2 and Assumptions 1-2, respectively. Assumption 3 puts a mild condition on the observation operator $\mathbf{H}_{t}$.
The following lemma provides a basic result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$ is an estimator of the Kalman Gain matrix assuming $\mathbf{R}_{t}$ is known. Closely related forms have appeared in the literature (Le Gland et al., 2009; Kwiatkowski and Mandel, 2015).
Lemma 3.1.
Suppose $\mathbf{R}_{t}$ is known and $\lambda_{\min}\left(\mathbf{R}_{t}\right)>0$, then
@@
Lemma 3.1 suggests that the upper bound of the estimation error in $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ mainly depends on the bound of $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|$, which is valid for both fixed and high dimensions. Note that under fixed dimension, the sample covariance is consistent with the underlying population covariance at the $\sqrt{n}$ rate, that is $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$ and thus $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|=O_{P}({n}^{-1/2})$. The corresponding results for the high dimensions are shown later.
Remark 1.
The last inequality in Lemma 3.1 is derived from $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|<1$ as $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}$ is the conditional variance of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=\boldsymbol{\Sigma}_{t}^{a}<\boldsymbol{\Sigma}_{t}^{f}$ as long as $\mathbf{x}_{t}$ is correlated with $\mathbf{y}_{1:t}$. Let us take a close look at $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$. For the extreme case of $\mathbf{R}_{t}=\mathbf{0}$, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be 0 if $\mathbf{H}_{t}$ is invertible. Intuitively, $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ would be close to 0 if $\mathbf{R}_{t}$ is close to $\mathbf{0}$. Suppose $p=q=1$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=(1+\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}/\sigma_{\mathbf{R}_{t}}^{2})^{-1}<1$. This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ decreases with the increase of $\mathbf{H}_{t}^{2}\sigma_{\boldsymbol{\Sigma}_{t}^{f}}^{2}$ or decrease of $\sigma_{\mathbf{R}_{t}}^{2}$. For $p,q>1$, if $\mathbf{H}_{t}=\mathbf{I}$, then $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|=\|\mathbf{I}-\boldsymbol{\Sigma}_{t}^{f}(\boldsymbol{\Sigma}_{t}^{f}+\mathbf{R}_{t})^{-1}\|=\varepsilon_{0}^{2}/(1+\varepsilon_{0}^{2})$ if we assume $\mathbf{R}_{t}=\varepsilon_{0}^{2}\boldsymbol{\Sigma}_{t}^{f}$. This means that $\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|$ will decrease with the decrease of $\varepsilon_{0}^{2}$, the ratio of $\mathbf{R}_{t}$ to $\boldsymbol{\Sigma}_{t}^{f}$.
In practice, the true observation error covariance matrix $\mathbf{R}_{t}$ may be unknown. Let $\hat{\mathbf{R}}_{t}$ be an estimator of $\mathbf{R}_{t}$ based on either knowledge or historical data. The following Corollary 3.1 provides a result that bounds $\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ for both fixed and high dimensional situations, where $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{\hat{R}}_{t}\big{)}^{-1}$ is an estimator of the Kalman Gain matrix when both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively.
Corollary 3.1.
If both the forecast and observation error covariance matrices $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{R}_{t}$ are estimated by $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ and $\hat{\mathbf{R}}_{t}$, respectively, then
$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|$ $\leq\big{(}\|\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t}\|\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{H}_{t}\|+\|\mathbf{K}_{t}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|\big{)}\|(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\hat{\mathbf{R}}_{t})^{-1}\|.$
If $\mathbf{R}_{t}$ is invertible and $\hat{\mathbf{R}}_{t}$ is invertible with probability approaching $1$, then
$\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|\leq{\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})}{\left\|\mathbf{H}_{t}\right\|}\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|+\lambda_{\min}^{-1}(\hat{\mathbf{R}}_{t})\lambda_{\min}^{-1}(\mathbf{R}_{t})\left\|\mathbf{H}_{t}\right\|\|\boldsymbol{\Sigma}_{t}^{f}\|\|\mathbf{R}_{t}-\hat{\mathbf{R}}_{t}\|.$
Compared with the upper bound in Lemma 3.1, there is an extra term related to the bound of the estimation error in the observation error covariance matrix $\mathbf{R}_{t}$. Most of the results established in the following of the paper are for the case of known $\mathbf{R}_{t}$ as is commonly assumed in the literature of the EnKF. Results for estimated $\mathbf{R}_{t}$ can be attained in the fashion shown in Corollary 3.1.
@@
Based on the result in Lemma 3.1 (or Corollary 3.1 for unknown $\mathbf{R}_{t}$), we can establish the convergence of the EnKF forecast and analysis states $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\mathbf{x}}_{t}^{a}$ to their oracle Kalman Filter counterparts $\mathbf{x}_{t}^{f}$ and $\mathbf{x}_{t}^{a}$ based on the true $\boldsymbol{\Sigma}_{t}^{f}$ and the true underlying evolution given in (2.4)-(2.5) by firstly establishing the convergence of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ to $\boldsymbol{\Sigma}_{t}^{f}$. For the fixed or low dimensional state vectors $\mathbf{x}_{t}$ where $p$ is of a smaller order of the ensemble size $n$ such that $p/n\to 0$, the sample covariance matrix is consistent such that $\|\hat{\boldsymbol{\Sigma}}_{t}^{f}-\boldsymbol{\Sigma}_{t}^{f}\|=O_{P}({n}^{-1/2})$.
However, remote sensing technology has been advancing the measurement of massive amounts of data sets for many geophysical processes, which produces high resolution observational data. Meanwhile, the increasing demands of high resolution forecast and dataset also make the spatiotemporal resolution of the state variables become much higher. Both make $\mathbf{x}_{t}$ high dimensional and bring high dimensional challenges on the validity of the EnKF. When the dimension $p$ of the state vector $\mathbf{x}_{t}$ is high so that $p\gg n$, the forecast ensemble sample covariance $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ in (2.8) is no longer consistent to the underlying covariance $\boldsymbol{\Sigma}_{t}^{f}$ (Bai and Yin, 1993), which is a major threat to the validity of the EnKF. The last two decades have witnessed rapid development in high dimensional statistical inference. Consistent high dimensional  covariance estimators have been proposed, which include the banding, tapering and thresholding estimators (Bickel and Levina, 2008b, a; Cai et al., 2010).
@@
Next, we present several consistent estimators of the high dimensional forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and their covariance classes before we evaluate the accuracy of the high dimensional assimilated state $\hat{\mathbf{x}}_{t}^{a}$ from the EnKF to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.
@@
3.1 Consistent estimation of high dimensional forecast error covariance matrix
The banding estimator (Bickel and Levina, 2008b) of the forecast error covariance matrix with bandwidth $k$ is
$B_{k}(\mathbf{S}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|i-j|\leq k)\big{]}_{p\times p},\,\,0\leq k<p,$
@@ 3.1 Consistent estimation of high dimensional forecast error covariance matrix
where $\mathbf{S}_{t}^{f}=[s_{tij}^{f}]_{p\times p}$ is the sample covariance of the EnKF forecast ensembles. The covariance matrix class suitable for the banding estimator is the following bandable covariance class, $\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{|i-j|>k}\left|\sigma_{ij}\right|\leq Ck^{-\alpha}\text{ for all }k>0\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}$ for some positive $\alpha$, $C$ and $\varepsilon_{0}$. This bandable pattern is generally consistent with the dependence between variables observed at sites $i$ and $j$ in geoscience studies, as the dependence becomes weaker as the distance between the two sites gets larger.
@@
The banding estimator administrates a hard threshold as it sets all matrix elements beyond the $k$-th sub-diagonals to zero. Cai et al. (2010) updated it with a tapering alteration which tapers down the level of the thresholding gradually
$\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)=\mathbf{S}_{t}^{f}\circ\mathbf{W}(k)=\big{[}w_{ij}(k)s_{tij}^{f}\big{]}_{p\times p},$
@@
where $\mathbf{W}(k)=(w_{ij}(k))_{p\times p}$, with $w_{ij}(k)=(2/k)\big{\{}(k-|i-j|)_{+}-(k/2-|i-j|)_{+}\big{\}},$ is the tapering matrix at a tapering width $k$ which is usually an even integer and $1\leq k\leq p$.
@@
Figure 1: Bandable $\mathcal{U}_{b}$ and circular bandable $\mathcal{U}_{bc}$ covariance matrices: the average energy function $h(k)$ in Panels (a) and (b) and their heatmaps in Panels (c) and (d), where $h(k)$ is the average of the squares of the $k^{th}$ sub-diagonal entries.
@@
In earth science, due to the largely spherical shape of the earth, for a study region that encompasses the entire earth or its poles, the state vectors are highly dependent for the variable components between the two ends of the state vector. Thus we propose a circular form of the bandable structure, $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\max_{j}\sum_{k_{1}<|i-j|<p-k_{2}}\left|\sigma_{ij}\right|\leq C(k_{1}+k_{2})^{-\alpha}\text{ for all }0<k_{1},k_{2}<p\text{ and }0<\varepsilon_{0}\leq\lambda_{\min}(\boldsymbol{\Sigma})\leq\lambda_{\max}(\boldsymbol{\Sigma})\leq 1/\varepsilon_{0}\big{\}}.$ Figure 1 depicts the $h(k)$ functions and heatmaps of two specific covariances that belong to the two bandable covariance classes $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$, where $h(k)$ is the average of the squares of the $k^{th}$ subdiagonal entries of a covariance defined by $h(k):=0.5{(p-k)^{-1}}\sum_{\left|l_{1}-l_{2}\right|=k}\sigma_{l_{1}l_{2}}^{2}=(p-k)^{-1}\sum_{l=1}^{p-k}\sigma_{l,l+k}^{2}.$ Figures 1 (a)-(b) show that the $h(k)$ functions of covariances in $\mathcal{U}_{b}$ and $\mathcal{U}_{bc}$ are $L$-type and $U$-type curves as the subdiagonal index $k$ increases, respectively. For covariances in the circular bandable class $\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, the proposed mid-banding forecast error covariance estimator with bandwidths $k_{1}$ and $k_{2}$ is
$B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})=[s_{tij}^{f}\mathbf{1}(|i-j|\leq k_{1}\,\text{or}\,|i-j|\geq(p-k_{2}))]_{p\times p},\,\,0\leq k_{1}<(p-k_{2})\leq p.$
@@
The bandable matrix class requires that the state vector $\mathbf{x}_{t}$ has an ordering (or a permutation) such that the correlation decays as two components of $\mathbf{x}_{t}$ are further apart. However, the presence of geo-spatial processes (weather front or hurricane) and geography (mountains or ocean) will interrupt the spatial dependence ordering prescribed by the banding covariance class. In such situations, the thresholding covariance class $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)$ suitable for the thresholding estimator (Bickel and Levina, 2008a) may be employed, where $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M,\varepsilon_{0}\right)=\big{\{}\boldsymbol{\Sigma}=(\sigma_{ij})_{p\times p}:\sigma_{ii}\leq M,\sum_{j=1}^{p}\left|\sigma_{ij}\right|^{\gamma}\leq c_{0}(p),\text{ for all }i\text{ and }\lambda_{\min}(\boldsymbol{\Sigma})\geq\varepsilon_{0}>0\big{\}},\text{ for }0\leq\gamma<1.$ The thresholding forecast error covariance matrix estimator with a threshold level $s$ is $T_{s}({\mathbf{S}}_{t}^{f})=\big{[}s_{tij}^{f}\mathbf{1}(|s_{tij}^{f}|\geq s)\big{]}_{p\times p}$.
@@
To apply the high dimensional estimators of the forecasting error covariance $\boldsymbol{\Sigma}_{t}^{f}$, we need the following distributional assumption.
Assumption 4.
The random vectors $\mathbf{x}_{t,i}^{f}=(x_{t,i1}^{f},\cdots,x_{t,ip}^{f})^{\mathrm{T}}$ are IID and each $x_{t,1j}^{f}\stackrel{{\scriptstyle d}}{{\sim}}F_{t,j}$ and $(x_{t,1j}^{f})^{2}\stackrel{{\scriptstyle d}}{{\sim}}G_{t,j}$, where $G_{t,j}(u)=F_{t,j}(\sqrt{u})-F_{t,j}(-\sqrt{u})$ for $u>0$. There exists a $\lambda_{0}>0$ such that $\max_{1\leq j\leq p}\mathbb{E}(e^{\lambda(x_{t,1j}^{f})^{2}})<\infty\textrm{ for }\lambda\in(-\lambda_{0},\lambda_{0}).$
Assumption 4 effectively assume $(x_{t,1j}^{f})^{2}$ are sub-exponentially distributed, which include the Sub-Gaussian random variables as a special case.
@@
Suppose the forecast ensemble $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are conditionally sub-exponentially distributed as in Assumption 4.
@@
    For the banding estimator $B_{k}(\mathbf{S}_{t}^{f})$ with $k\asymp(n^{-1}\log p)^{-1/(2\alpha+2)}$, then uniformly for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$